OpenStudy (jkbo):

Please helpWithout graphing, classify the following system as independent, dependent, or inconsistent. -21x-3y=-24 y=-7x+7 a.independent b.dependent c.inconsistent

OpenStudy (anonymous):

from the first equation you can take out 3 as a common factor and reduce it to y = -7x+8. Combined with y = -7x +7 what kind of system does it form?

OpenStudy (anonymous):



OpenStudy (anonymous):

parallel lines with a slope of -7

OpenStudy (jkbo):

ok so its indepent

OpenStudy (anonymous):

for a system of linear equations a1x + b1y + c1 = 0 and a2x + b2y + c2 = 0 if a1/a2 not equal to b1/b2 - the two lines meet at a point and have unique solution - so consistent and independent if a1/a2 = b1/b2 = c1/c2 - the two lines coincide and so infinite solutions - so consistent and dependent if a1/a2 = b1/b2 not equal to c1/c2 - the two lines are parallel and hence no solution - so inconsistent

OpenStudy (jkbo):

o ok thank you very much

OpenStudy (anonymous):

Your given equations can be written as -21x-3y + 24 = 0 and 7x + y - 7 = 0 a1/a2 = -21/7 = -3/1 b1/b2 = -3/1 c1/c2 = 24/-7 = -24/7 so a1/a2 =b1/b2 not equal to c1/c2 so two lines are parallel--hence no solution--hence inconsistent

OpenStudy (jkbo):

@Harkirat Can you help me with one more question. Graph the system of inequalities, Which two quadrants does the solution lie in? y > x+2 y<-x+1 A.Quadrants 2 and 3 B.Quadrants 3 and 4 C.Quadrants 2 and 4 D.Quadrants 1 and 2

OpenStudy (anonymous):

i think B.Quadrants 3 and 4

OpenStudy (anonymous):

You will have to plot the graphs to see how it turns out....

OpenStudy (jkbo):

@Harkirat yeah I need help, how do you know what quadrants because the lines go in different directions